OpenStudy (anonymous):

Translate the word phrase into a math expression. the product of twenty-four and the difference between ten and two A. 24 ÷ (10 – 2) B. 24 • 10 – 2 C. 24 ÷ (10 + 2) D. 24 • (10 – 2)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

you have two things here and their product, so its x * y, where x is first thing, and y is second thing, which leads to answer you picked :)

OpenStudy (anonymous):

it's true :)

OpenStudy (anonymous):

product is x right

OpenStudy (anonymous):

so D is right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x = 24 and y = 10-2, so x*y = 24*(10-2)